# Patterns of imbalance states between sub-brain regimes during development in the resting state

### v1

@@ Introduction
A closer examination of anticorrelated subnetworks across development reveals a dynamic trajectory of functional segregation and integration that closely parallels cognitive maturation. Functional segregation refers to the extent to which regions within a given brain network are connected to each other, yet relatively connected to other networks. This supports specialized processing within domains such as attention, memory, and sensory input. In contrast, functional integration denotes the coordination and communication between distinct brain networks. This allows for flexible, goal-directed behavior and cross-domain cognitive control sporns2013network ; wig2017segregated . Optimal cognitive functioning requires balancing these two principles: segregation ensures efficiency and specialization, and integration enables adaptive behavior by flexibly combining information across systems. In childhood, clear anticorrelations are already present, particularly between the default mode network (DMN) and the dorsal attention network (DAN), as well as between the DMN and salience/ventral attention network (VAN). These early negative functional connections reflect the developing ability to differentiate between internally oriented thought (e.g., autobiographical memory, mind-wandering) and externally driven attentional systems fransson2007resting ; gao2017functional . The emergence of these patterns as early as age 2–4 suggests that functional segregation begins early, providing a neural foundation for attentional control and self-regulation in middle childhood fair2008maturing . In adolescence, however, this architecture undergoes a period of reorganization. While core anticorrelations—such as DMN–DAN—remain detectable, they are reduced in strength, consistent with a transient weakening of network boundaries due to synaptic pruning, white matter remodeling, and dynamic reconfiguration baum2017modular ; marek2015contribution . This “dip” in functional segregation during adolescence may support increased flexibility as the brain reorganizes and consolidates more efficient long-range connections. Importantly, these changes parallel behavioral shifts such as increased cognitive flexibility, emerging abstract reasoning, and adaptive emotional regulation luna2010has . By adults, the brain reaches a period of peak network segregation, where within-network connectivity is maximized and between-network anticorrelations are strongest chan2014decreased ; nomi2017chronnectomic . Notably, strong anticorrelations emerge not only between the DMN and DAN, but also between the DMN and the frontoparietal control network (FPN), a configuration that supports efficient switching between internally generated thought and goal-directed behavior spreng2013intrinsic . Other studies show that additional anticorrelations involving limbic, visual, somatomotor, and salience networks also become more prominent, suggesting greater modularity and cross-network coordination betzel2014changes . These changes underpin adult-level cognitive functions such as sustained attention, working memory, and top-down executive control andrews2014default ; satterthwaite2013heterogeneous .
Figure 6: Anti correlation between subnetworks for different age groups: a) Children, b) Adolescents, c) Adults.
@@
SalVentAttn Cont 19.0 0.039
Cont Default 17.0 0.035
DorsAttn Cont 13.0 0.026
DorsAttn DorsAttn 11.0 0.022
SomMot Default 10.0 0.020
SalVentAttn SomMot 9.0 0.018
Visual Default 5.0 0.010
Visual Default 3.0 0.006
Visual Default 2.0 0.004
Cont Cont 1.0 0.002
SomMot Limbic 1.0 0.002
Default Limbic 1.0 0.002
SomMot Cont 1.0 0.002
SomMot DorsAttn 1.0 0.002
Table 2: Subnetworks connection weights for Children
@@
SalVentAttn Cont 19.0 0.090
DorsAttn Cont 15.0 0.071
DorsAttn SalVentAttn 10.0 0.048
Visual DorsAttn 9.0 0.043
Default Limbic 7.0 0.033
Visual Default 6.0 0.029
DorsAttn Limbic 5.0 0.024
Default SomMot 4.0 0.019
Visual Cont 3.0 0.014
DorsAttn SomMot 3.0 0.014
DorsAttn DorsAttn 3.0 0.014
SalVentAttn Limbic 2.0 0.010
SalVentAttn SomMot 2.0 0.010
Cont SomMot 1.0 0.005
Table 3: Subnetworks connection weights for Adolescents
@@
Limbic Default 82.0 0.102
SalVentAttn Cont 82.0 0.102
DorsAttn Cont 58.0 0.072
Default Default 42.0 0.052
SomMot DorsAttn 41.0 0.051
SomMot SalVentAttn 37.0 0.046
DorsAttn Limbic 34.0 0.042
Visual Default 28.0 0.035
DorsAttn SalVentAttn 23.0 0.029
Visual SalVentAttn 18.0 0.022
SalVentAttn Limbic 16.0 0.020
DorsAttn Visual 14.0 0.017
Visual Cont 10.0 0.012
SomMot Cont 9.0 0.011
SomMot SomMot 8.0 0.010
SomMot Limbic 6.0 0.007
Visual Limbic 6.0 0.007
Cont Cont 3.0 0.004
Limbic Cont 2.0 0.002
Table 4: Subnetworks connection weights for Adults

### v2

@@ Introduction
A closer examination of anticorrelated subnetworks across development reveals a dynamic trajectory of functional segregation and integration that closely parallels cognitive maturation. Functional segregation refers to the extent to which regions within a given brain network are connected to each other, yet relatively connected to other networks. This supports specialized processing within domains such as attention, memory, and sensory input. In contrast, functional integration denotes the coordination and communication between distinct brain networks. This allows for flexible, goal-directed behavior and cross-domain cognitive control sporns2013network ; wig2017segregated . Optimal cognitive functioning requires balancing these two principles: segregation ensures efficiency and specialization, and integration enables adaptive behavior by flexibly combining information across systems. In childhood, clear anticorrelations are already present, particularly between the Default Mode Network (DMN) and the Dorsal Attention Network (DAN), as well as between the DMN and the Salience/Ventral Attention Network (VAN). These early negative functional connections reflect the developing ability to differentiate between internally oriented thought (e.g., autobiographical memory, mind-wandering) and externally driven attentional systems fransson2007resting ; gao2017functional . The emergence of these patterns as early as age 2–4 suggests that functional segregation begins early, providing a neural foundation for attentional control and self-regulation in middle childhood fair2008maturing . In adolescence, however, this architecture undergoes a period of reorganization. While core anticorrelations—such as DMN–DAN—remain detectable, they are reduced in strength, consistent with a transient weakening of network boundaries due to synaptic pruning, white matter remodeling, and dynamic reconfiguration baum2017modular ; marek2015contribution . This “dip” in functional segregation during adolescence may support increased flexibility as the brain reorganizes and consolidates more efficient long-range connections. Importantly, these changes parallel behavioral shifts such as increased cognitive flexibility, emerging abstract reasoning, and adaptive emotional regulation luna2010has . By adults, the brain reaches a period of peak network segregation, where within-network connectivity is maximized and between-network anticorrelations are strongest chan2014decreased ; nomi2017chronnectomic . Notably, strong anticorrelations emerge not only between the DMN and DAN, but also between the DMN and the Frontoparietal Control Network (FPN), a configuration that supports efficient switching between internally generated thought and goal-directed behavior spreng2013intrinsic . Other studies show that additional anticorrelations involving limbic, visual, somatomotor, and salience networks also become more prominent, suggesting greater modularity and cross-network coordination betzel2014changes . These changes underpin adult-level cognitive functions such as sustained attention, working memory, and top-down executive control andrews2014default ; satterthwaite2013heterogeneous .
(a) Children
(b) Adolescents
(c) Adults
Figure 6: (a–c) Networks of anticorrelations between functional subnetworks in children, adolescents, and adults. Nodes represent subnetworks, while edge weights indicate the strength of negative connectivity. Results highlight the developmental trajectory from localized anticorrelations to strongly segregated and integrated subnetworks in adulthood.
@@
Table 1: Major anticorrelated subnetworks identified from $\Delta{{}^{*}}_{---}$ in childhood. The listed weights reflect the strength of anticorrelation between subnetworks, with proportions showing their relative dominance within the network.
@@
Table 2: Major anticorrelated subnetworks identified from $\Delta{{}^{*}}_{---}$ in adolescence. The listed weights reflect the strength of anticorrelation between subnetworks, with proportions showing their relative dominance within the network.
@@
Table 3: Major anticorrelated subnetworks identified from $\Delta{{}^{*}}_{---}$ in adulthood. The listed weights reflect the strength of anticorrelation between subnetworks, with proportions showing their relative dominance within the network.